### v1

@@ 3 The Regulatory Framework Based on the DSA and AIA on Third-Party Audits and Data Access
The goal of this chapter is to understand how the DSA and AIA regulate third-party audits and third-party data and model access. Therefore we will analyze the the DSA on audits and data access first. Subsequently, we will do the same for the current draft of the AIA. It should be noted that the following explanations refer to the draft of the European Parliament’s proposal.
@@ 3.1 Digital Services Act
In order to ensure compliance with the DSA, a “novel institutional ecosystem” [46] will be established that involves independent third parties in an additional oversight role. [^5] Cf. Recital No. 92 DSA. [27] Therefore, the officially promulgated version of the DSA asserts that VLOPs and VLOSEs are legally required to undergo an independent audit at least once a year (“independent audit”) [^6] Art. 37 (1) DSA. The purpose of auditing is to determine whether the providers are following the obligations standardized by the DSA. [^7] It means the obligations set out in Chapter III (Art. 11-48 DSA), in particular the obligations arising from codes of conduct referred to in Art. 45 and 46 DSA and the crisis protocols referred to in Art. 48 DSA. This is an important regulatory requirement, as it increases providers’ accountability and enables regulators and the public to understand and regulate how VLOPs and VLOSEs moderate content. The Commission is empowered to implement the requirements of Art. 37 DSA by legislative act. It has published a first draft of a regulation to specify the requirements for independent audits (hereinafter referred to as the Delegated Regulation on independent audits) [24]. The introduction of such guidelines is welcome as it will provide legal certainty for both audit firms and VLOPs and VLOSEs by providing guidance that can improve transparency and accountability. An important audit tool is the preparation of an audit report, which should comply with a comprehensive set of obligations. The audit firm must describe the specific elements audited and the methodology used. [^8] Art. 37 (4)(2d) DSA. It must also provide a description and summary of the main audit findings. [^9] Art. 37 (4)(2e) DSA. In addition, opinion must be given as to whether the audited provider has complied with the obligations and commitments referred to above. [^10] Art. 37 (4)(2g) DSA. The audit opinion shall state the result, which can be “positive”111111“Positive” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has complied with an audited obligation or commitment, Art. 8 (1)(a) Delegated Regulation on Independent Audits [24]., “positive with comments”121212“Positive with comments” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has complied with an audited obligation or commitment, Art. 8 (1)(b) Delegated Regulation on Independent Audits. or “negative”131313“Negative” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has not complied with an audited obligation or commitment, Art. 8 (1)(c) Delegated Regulation on Independent Audits. In the case of a negative remark, the audit firm must make operational recommendations to be implemented within a recommended period of time to achieve compliance. [^14] Art. 37 (4)(2h) DSA If the audit company cannot conduct the auditing at all or can only do so partially, this should be clarified. [^15] Art. 37 (5) DSA. In such cases the Delegated Regulation on independent audits could specify how the systematic assessment of risks must be conducted. It should also specify the factors to be taken into account in the risk analysis. [^16] Cf. Art. 9 (4) Delegated Regulation on independent audits. According to the current draft the inherent risks171717“Inherent risks” means the risk of non-compliance intrinsically related to the nature, the activity and the use of the audited service, as well as the context in which it operates, and the risk of non-compliance related to the nature of the audited obligation or commitment, Art. 2 (10) Delegated Regulation on Independent Audits., the control risk181818“Control risks” means the risk that a misstatement is not prevented, detected and corrected in a timely manner by means of the audited provider’s internal controls, Art. 2 (11) Delegated Regulation on Independent Audits, and the detection risks191919“Detection risk” means the threshold beyond which deviations or misstatements by the audited provider, individually or aggregated, would reasonably affect the audit findings, conclusions and opinions, Art. 2 (12) Delegated Regulation on Independent Audits. should accounted for in the audit companies’ analysis. [^20] Art. 9 (3) Delegated Regulation on Independent Audits. Providers should provide the necessary cooperation and assistance to organizations conducting audits under this obligation so that they can undertake their work in an effective, efficient, and timely manner. [^21] Art. 37 (2)(1) DSA. They must, for example, grant access to all relevant data and premises. This includes the disclosure of data relating to algorithmic systems and answering written or oral questions. [^22] Cf. Recital No. 92 DSA. In addition to this, the Delegated Regulation on independent audits could regulate which information the providers have to provide to the audit organisations, Art. 5 (1) Delegated Regulation on Independent Audits. The provider and the audit organisation must necessarily agree on conditions - i.e. meaning duties and obligations - that are listed in the current draft to carry out the audit, Art. 7 (1)(a) Delegated Regulation on Independent Audits. The platform provider must then consider improvement suggestions and take action to enforce the recommendations within one month. [^23] Art. 37 (6)(2) DSA. If the operational recommendations are not implemented, the VLOP or VLOSE provider must justify and specify alternative measures that will be taken to fix the identified cases of non-compliance. [^24] Art. 37 (6)(3) DSA.
@@ 3.1.2 The Requirements for Audit Firms under the DSA
Art. 40 DSA regulates access to the data of VLOPs and VLOSEs for “vetted researchers”.The norm contains not only the obligation of providers to grant access, but also the right of researchers to access.  [^28] E. g. [38] and [91] Researchers must undertake to make their research results publicly available free of charge  [^29] Art. 40 (8)(1g) DSA.. The results can thus be made available to the authorities, the Commission and the public. Classification as an approved researcher is subject to strict conditions. In particular, researchers must be affiliated to a research institution  [^30] Art. 40 (8)(1a) DSA.. It means a university, including its libraries, a research institute or any other entity, the primary goal of which is to conduct scientific research or to carry out educational activities involving also the conduct of scientific research.  [^31] Art. 2 (1) Directive (EU) 2019/790. Like the audit firms, the researchers and the institutions to which they belong must be independent of commercial interests  [^32] Art. 40 (8)(1b) DSA.. Another condition for access is that the data collection is necessary and proportionate for the creation of the data work.  [^33] Art. 40 (8)(1e) DSA. Cf. Recital 97. The classification as vetted researcher is granted to the researcher only in relation to the intended research project. The research must relate to systemic risks and contribute to their identification and understanding. [^34] Art. 40 (4) and (8f) DSA.
@@ 3.2 The Artificial Intelligence Act
The AIA [22, 26] aims to regulate and harmonise the marketing, commissioning and use of AI systems within the European Union. To this end, safe and trustworthy AI is to be created by observing the fundamental values protected in the European Union, that is material and ethical values, during development and use. For this reason, companies [^35] The regulation is addressed to all companies that place AI systems on the market or put them into operation in the European Union, Art. 2 (1) AIA [22, 26]. In practice, this will include software developers, AI importers and distributors, but also, for example, government authorities. should be subject to a far-reaching catalogue of obligations and ensure quality assurance throughout the entire AI value chain. In contrast to the Commission, the European Parliament’s draft also takes account of foundation models, which would be subject to specific regulatory requirements [96].
@@ 3.2.1 The Regulations on the Conformity Assessment Procedure
Basically, in the context of conformity assessment procedures, the AIA focuses primarily on internal control by the first parties. [^36] Annex VI AIA. The AIA opens up the possibility of having such procedures performed by external companies only for so-called high-risk systems [^37] Cf. Art. 6 AIA and Annex III for a description of which AI systems could be high-risk system.. [^38] Art. 43 (1)(1b) AIA. However, whether a provider of an AI system uses a notified body depends on the nature of the high-risk system. According to the current draft, it remains optional if, for example, biometric and biometrics-based systems are used, which are partly subject to harmonised standards that cover the relevant security requirements for the AI system in the AIA and common specifications exist in accordance with Art. 41 AIA. [^39] Art. 43 (1)(subpar. 1) AIA: “[…] the provider shall opt for one of the following procedures:”. On the other hand, the conformity assessment procedure by external bodies is mandatory, for example, when harmonised standards are only partially applicable to the AI system or if the provider considers an external audit to be necessary. [^40] Cf. Art. 43 (1)(subpar. 2) AIA. An audit should also be carried out if the AI system is already subject to a conformity assessment procedure, but there has been a significant change to the system, or the change is as a result of the system’s learning process. [^41] Cf. Art. 43 (4) AIA. In the other cases, [^42] This refers to high-risk systems in the sense of Annex III Nr. 2-8 AIA. the AI systems are only subject to internal control. They are not subject to conformity assessment procedure within the meaning of the AIA [^43] Cf. Art. 43 (2)(1) AIA. or a conformity assessment procedure according to other regulations under European law. [^44] Art. 43 (3)(subpar. 1)(1) AIA. In the latter case, however, the provider of the AI system must take into account some of the procedural requirements of the AIA. [^45] Art. 43 (3)(subpar. 1)(2) AIA. The requirements for the conformity assessment procedure are set out in Annex VII. The procedure involving a notified body should be conducted in a way that preserves the confidentiality of the information provided. [^46] Cf. Art. 33 (6) and 70 (1) AIA. It is composed of the audit and monitoring of the respective approved quality management system, [^47] Annex VII No. 3, No. 5 AIA as well as the control of the technical documentation obligations [^48] Annex VII No. 4 AIA. imposed on the providers. Furthermore, the audit firm may be granted access to the training models and trained models of the AI system, including its relevant parameters upon request, and they should be able to test the AI system independently. [^49] Annex VII No. 4.4 and No. 4.5 AIA Provided that the notified body can confirm the conformity of the high-risk system, a certificate — similar to the audit opinion under the DSA — should be issued. [^50] Art. 44 (1) and (2) AIA. That certificate would have a duration of up to five years or could be revoked or limited if the AI system later fails to meet the requirements of the AIA. [^51] Art. 44 (3) AIA. The providers of the AI system should support the audit organizations in their investigations by, for example, providing the technical documentation that makes available all necessary information to assess whether the AI system complies with these requirements. [^52] Art. 11 (1) AIA. Non-compliance with the support services is subject to a fine of up to EUR 5 million or 1 per cent of total annual worldwide turnover, Art. 71 (5) AIA. They have to grant the audit company unrestricted access to the training and test data sets that the provider used. This also applies to application programming interfaces and other means and tools suitable for remote access. [^53] Annex VII No. 4.3 AIA. In order to adapt the regulatory framework, the power to adopt acts amending the provisions on conformity assessment procedures should be delegated to the Commission. [^54] Recital No 85 (1) AIA.
@@ 3.2.2 The Requirements for Audit Firms under the AIA
So-called “notified bodies” should perform the conformity assessment procedures. [^55] Art. 33 (1) AIA. A “notified body” is defined in the AIA as a conformity assessment body notified in accordance with the AIA and other relevant Union harmonisation legislation. [^56] Art. 3 (1)(22) AIA. The competent authority [^57] This is the national authority of the respective Member State to which the monitoring of compliance with the AI Act has been assigned. According to Art. 3 (1)(19) AIA, a designated authority is the national authority responsible for establishing and carrying out the necessary procedures for the assessment, designation, and notification of conformity assessment bodies and for their monitoring. Among the tasks of the notifying authority, cf. Art. 59 AIA. then notifies them, provided that they fulfil the specific requirements in the AIA. [^58] Art. 31 (1), 32 (1) AIA. The procedure for designating qualified bodies specifies requirements for (professional) qualifications as well as their independence and objectivity. As in the DSA, the AIA also clarifies that the notified bodies should be independent of the provider of the high-risk system. [^59] Art. 33 (4) and (5) AIA. Cf. also the European Parliament’s clarification that the assessment may only be carried out by staff who have neither worked for the high-risk scheme provider nor been employed in a related legal entity for twelve months prior to the assessment, Art. 33 (4a) AIA. In this context, the independence of the notified body extends to the provider of an assessed AI system as well as to all other actors with an economic interest in thas AI system. [^60] Cf. Art. 33 (4) AIA. In order to ensure professional competence, they must meet the specific organisational, quality management, resourcing, and procedural requirements necessary for the conformity assessment of AI systems. [^61] Cf. Art. 33 (2), (3), (10) AIA. To this end, the notified bodies should have procedures guiding their activities which take due account of the size of an undertaking, the sector in which it operates, its structure and the degree of complexity of the AI system concerned. [^62] Cf. Art. 33 (9) AIA. In principle, the provider of the AI system can choose between the notified bodies, unless the AI system is put into operation by a law enforcement, immigration or asylum authority or by EU institutions, bodies or agencies. [^63] Cf. Art. 43 (1)(subpar. 3) AIA. Then the market surveillance authority mentioned in Art. 63 (5) or (6) AIA takes over the function of the notified body. In contrast to the DSA, the AIA ensures close cooperation with the notifying authority. The state participation control is enforced by the certification of the bodies to be inspected. Furthermore, the certification bodies are accountable to the notifying authorities. In order to qualify as a notified body in the sense of the AIA, the potential audit company has to fulfil a number of requirements: Extensive information must be added to the application for notification, such as the conformity assessment activities, the conformity assessment modules and the AI technologies for which the conformity assessment body claims competence, and – if available – an accreditation certificate. [^64] The accreditation certificate is issued by a national accreditation body and certifies that the conformity assessment body complies with the requirements of the AIA, Art. 31 (2) AIA. In the absence of a corresponding accreditation certificate, the necessary evidence for the verification, recognition and periodic monitoring of compliance with the requirements of Art. 33 AIA shall be sent to the designated authority in accordance with AIA[56].

### v3

@@ 3 The Regulatory Framework Based on the DSA and AIA on Third-Party Audits and Data Access
The objective of this chapter is to comprehend the regulations of third-party audits and third-party data and model access by the DSA and AIA. We will begin by examining the DSA’s provisions on audits and data access (section 3.1). Then, we will do the same for the AIA based on the Corrigendum made public by the European Parliament (section 3.2).
@@ 3.1 Digital Services Act
In order to ensure compliance with the DSA, a “novel institutional ecosystem” [50] will be established that involves independent third parties in an additional oversight role. [^6] Cf. Recital No. 92 DSA. [31] Therefore, the officially published version of the DSA asserts that VLOPs and VLOSEs are legally required to undergo an independent audit at least once a year (“independent audit”) [^7] Art. 37 (1) DSA. The purpose of auditing is to determine whether the providers are following the obligations standardized by the DSA. [^8] It means the obligations set out in Chapter III (Art. 11-48 DSA), in particular the obligations arising from codes of conduct referred to in Art. 45 and 46 DSA and the crisis protocols referred to in Art. 48 DSA. This is an important regulatory requirement, as it increases providers’ accountability and enables regulators and the public to understand and regulate how VLOPs and VLOSEs moderate content. The Commission is empowered to implement the requirements of Art. 37 DSA by legislative act. It has published a first draft of a regulation to specify the requirements for independent audits (hereinafter referred to as the Delegated Regulation on independent audits) [27]. The introduction of such guidelines is welcome as it will provide legal certainty for both audit firms and VLOPs and VLOSEs by providing guidance that can improve transparency and accountability. An important audit tool is the preparation of an audit report, which should comply with a comprehensive set of obligations. The audit firm must describe the specific elements audited and the methodology used. [^9] Art. 37 (4)(2d) DSA. It must also provide a description and summary of the main audit findings. [^10] Art. 37 (4)(2e) DSA. In addition, an outcome statement must be given as to whether the audited provider has complied with the obligations and commitments referred to above. [^11] Art. 37 (4)(2g) DSA. The audit statement shall state the result, which can be “positive” 121212“Positive” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has complied with an audited obligation or commitment, Art. 8 (1)(a) Delegated Regulation on Independent Audits [27]., “positive with comments”131313“Positive with comments” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has complied with an audited obligation or commitment, Art. 8 (1)(b) Delegated Regulation on Independent Audits. or “negative”141414“Negative” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has not complied with an audited obligation or commitment, Art. 8 (1)(c) Delegated Regulation on Independent Audits. In the case of a negative remark, the audit firm must make operational recommendations to be implemented within a recommended period of time to achieve compliance. [^15] Art. 37 (4)(2h) DSA If the audit company cannot conduct the auditing at all or can only do so partially, this should be clarified. [^16] Art. 37 (5) DSA. In such cases the Delegated Regulation on independent audits could specify how the systematic assessment of risks must be conducted. It should also specify the factors to be taken into account in the risk analysis. [^17] Cf. Art. 9 (4) Delegated Regulation on independent audits. According to the current draft, the inherent risks181818“Inherent risks” means the risk of non-compliance intrinsically related to the nature, the activity and the use of the audited service, as well as the context in which it operates, and the risk of non-compliance related to the nature of the audited obligation or commitment, Art. 2 (10) Delegated Regulation on Independent Audits., the control risk191919“Control risks” means the risk that a misstatement is not prevented, detected and corrected in a timely manner by means of the audited provider’s internal controls, Art. 2 (11) Delegated Regulation on Independent Audits, and the detection risks202020“Detection risk” means the threshold beyond which deviations or misstatements by the audited provider, individually or aggregated, would reasonably affect the audit findings, conclusions and statements, Art. 2 (12) Delegated Regulation on Independent Audits. should be accounted for in the audit companies’ analysis. [^21] Art. 9 (3) Delegated Regulation on Independent Audits. Providers should provide the necessary cooperation and assistance to organizations conducting audits under this obligation so that they can undertake their work in an effective, efficient, and timely manner. [^22] Art. 37 (2)(1) DSA. They must, for example, grant access to all relevant data and premises. This includes the disclosure of data relating to algorithmic systems and answering written or oral questions. [^23] Cf. Recital No. 92 DSA. In addition to this, the Delegated Regulation on independent audits could regulate which information the providers have to provide to the audit organisations, Art. 5 (1) Delegated Regulation on Independent Audits. The provider and the audit organisation must necessarily agree on conditions - i.e. meaning duties and obligations - that are listed in the current draft to carry out the audit, Art. 7 (1)(a) Delegated Regulation on Independent Audits. The platform provider must then consider improvement suggestions and take action to enforce the recommendations within one month. [^24] Art. 37 (6)(2) DSA. If the operational recommendations are not implemented, the VLOP or VLOSE provider must justify and specify alternative measures that will be taken to fix the identified cases of non-compliance. [^25] Art. 37 (6)(3) DSA.
@@ 3.1.2 The Requirements for Audit Firms under the DSA
Art. 40 DSA regulates access to the data of VLOPs and VLOSEs for “vetted researchers.” The norm contains not only the obligation of providers to grant access, but also the right of researchers to access.  [^29] E. g. [42] and [98] Researchers must make their research results publicly available free of charge  [^30] Art. 40 (8)(1g) DSA.. The results can thus be made available to the authorities, the Commission and the public. Classification as an approved researcher is subject to strict conditions. In particular, researchers must be affiliated to a research institution  [^31] Art. 40 (8)(1a) DSA.. It means a university, including its libraries, a research institute or any other entity, the primary goal of which is to conduct scientific research or to carry out educational activities involving also the conduct of scientific research.  [^32] Art. 2 (1) Directive (EU) 2019/790. Like the audit firms, the researchers and the institutions to which they belong must be independent of commercial interests  [^33] Art. 40 (8)(1b) DSA.. Another condition for access is that the data collection is necessary and proportionate for the creation of the data work.  [^34] Art. 40 (8)(1e) DSA. Cf. Recital 97. The classification as vetted researcher is granted to the researcher only in relation to the intended research project. The research must relate to systemic risks and contribute to their identification and understanding. [^35] Art. 40 (4) and (8f) DSA.
@@ 3.2 The Artificial Intelligence Act
The AIA [25, 29] aims to regulate and harmonise the marketing, commissioning and use of AI systems within the European Union. To this end, safe and trustworthy AI is to be created by observing the fundamental values protected in the European Union, that is material and ethical values, during development and use. For this reason, companies [^36] The regulation is addressed to all companies that place AI systems on the market or put them into operation in the European Union, cf. Art. 2 (1) AIA [25, 29]. In practice, this will include software developers, AI importers and distributors, but also, for example, government authorities. should be subject to a far-reaching catalogue of obligations and ensure quality assurance throughout the entire AI value chain.
@@ 3.2.1 The Regulations on the Conformity Assessment Procedure
The AIA only regulates conformity assessment requirements for high-risk AI systems. For this purpose, the AIA defines “conformity assessment” as “the process of demonstrating whether the requirements […] to a high-risk AI system have been fulfilled.” [^37] Art. 3 (20) AIA. The aim is thus to promote the accountability of AI providers. On the one hand, a so-called “embedded” AI system [^38] Art. 6 (1) AIA and Annex I. These are, for example, products for machines, medical devices, automobiles and aircraft, Rec. (50) AIA. is subject to conformity assessment by third parties. AI systems are already subject to other product safety regulations; the conformity assessments are carried out in accordance with these different European regulations.  [^39] Art. 43 (3)(1) AIA. Cf. Rec. (49),(51) AIA. In this case, however, the provider of the AI system must take into account some of the procedural requirements of the AIA. [^40] Art. 43 (3) AIA. As these AI systems are already subject to detailed safety requirements, which are assessed under a public authorisation procedure, the AIA is intended to complement these requirements. [^41] The regulation of the systems is based on the legal provisions of the New Legislative Framework. This means that the product safety requirements for a system should be based on an EU Directive, with the standardisation organisations CEN, CENELEC and ETSI providing more specific details. For so-called stand-alone AI systems [^42] Art. 6 (2) AIA and Annex III., the conformity assessment must generally be carried out internally, i.e. by the providers themselves. With regard to these systems, it is explicitly stated that the involvement of a so-called notified body is not necessary for the assessment. [^43] Art. 43 (2) AIA. Due to the complexity of these AI systems and the involved risks, the legislator assumes that the suppliers are best placed to evaluate them. This also means that audits that potentially are part of such a conformity assessment can be conducted internally, thus implying first-party auditing. However, this does not apply to AI systems that are used as ”remote biometric identification systems” (Annex III No. 1(a) AIA.). For these, it is optional whether the conformity assessment is carried out by the provider himself or by a notified body. [^44] Art. 43 (1) AIA. Furthermore, the conformity assessment by external bodies is mandatory for a few cases, e.g. when the AI system is put into operation by law enforcement or immigration authorities or by institutions, bodies, offices or agencies of the Union. [^45] Cf. Art. 43(1), (2) AIA. For general purpose AI systems, cf. Rec. (161) AIA. Another requirement for specific ’high-risk’ AI systems  [^46] Cf. Art 27 (1) AIA. Mandatory for high-risk AI system as described in Article 6(2), with the exception of those intended for use in the areas listed in point 2 of Annex III, deployers that are bodies governed by public law, or are private entities providing public services, and deployers of high-risk AI systems referred to in points 5 (b) and (c) of Annex III. is the execution of a fundamental rights impact assessment, however, also conducted by the deployers similar and an extension to data protection impact assessments [91].
@@ 3.2.2 The Requirements for Audit Firms under the AIA
A ’notified body’ is defined in the AIA as “a conformity assessment body notified in accordance with this Regulation and other relevant Union harmonisation legislation.” [^47] Art. 3 (1)(22) AIA. For the term ’conformity assessment body’, see Art. 3 (1)(21). Cf. for the term ’notifying authority’ Art. 3 (19) and Art. 28 (1) AIA. For the duties of the notifying authority, cf. Art. 3 (19) and 28-30 AIA. The procedure for their designation sets out requirements for their (professional) qualifications as well as their independence and objectivity. As in the DSA, the AIA also clarifies that notified bodies should be independent of the provider of the high-risk system. This includes both economic aspects, e.g. no competitive relationship between the provider of the AI system and the notified body, and personnel aspects, e.g. no connection between the management and the AI development project. [^48] Art. 31(4) and (5) AIA. In order to ensure technical competence, they have to note specific organisational, quality management, personnel and procedural requirements. [^49] Art. 31(2), (3), (10) and Rec. (126) AIA. To this end, notified bodies should have procedures in place to guide their activities, taking into account the size of an undertaking, the sector in which it operates its structure, and the degree of complexity of the AI system concerned. [^50] Cf. 31(8) AIA. In addition, it should be noted that the conformity assessment procedure for AI systems, that will be used by certain authorities and related bodies (e.g. law enforcement, immigration or asylum authorities or EU bodies) is carried out by the market surveillance authority. [^51] Cf. Art. 43(1) AIA. Unlike the DSA, the AIA ensures the cooperation between the developer, the notified body, and the notifying authority. The authority is enforced to certify the bodies. In addition, the certificated bodies are accountable to the notifying authorities. To qualify as a notified body under the AIA, the potential audit firm must fulfill several requirements: The notification application must be accompanied by comprehensive information, such as the conformity assessment activities, conformity assessment modules, and AI technologies for which the conformity assessment body claims competence, as well as - if available - an accreditation certificate. [^52] Cf. Art. 29 (2)(3) AIA. The provider shall ensure that the conformity assessment process takes place. [^53] Cf. Art. 6(f) AIA. An audit should also be carried out if the AI system has already undergone a conformity assessment procedure, but a significant change has been made to the system. [^54] Cf. Art. 43(4) AIA, for the definition in Art. 3(15) AIA. In any case, they are obliged to document the assessment in order to provide evidence that the conformity assessment has been carried out properly. [^55] Art. 18 (1)(g), Art. 22 (3)(b) AIA. The requirements for the conformity assessments by internal controls are set out in Annex VI. It consists of the establishment and monitoring of the relevant quality management system, [^56] Annex VI No. 2, No. 4 AIA as well as the control of the obligations imposed on the suppliers regarding the technical documentation. [^57] Annex VI No. 3 AIA The providers of the AI system should assist the testing organisations in their investigations, e.g. by making available the technical documentation that provides all the information necessary to assess whether the AI system meets these requirements. [^58] Cf. Rec. (86) AIA. Providers must give them full access to the training and test data sets used. This includes application programming interfaces and other means and tools suitable for remote access. [^59] Annex VII No. 4.3 AIA. For example, the audit firm may, upon request, be given access to the AI system’s training models and trained models, including the relevant parameters, and should be able to test the AI system independently. [^60] Annex VII No. 4.4 and 4.5 AIA. For this reason, it is appropriate for the notified body to maintain the confidentiality of the information provided. [^61] Cf. Art. 31(7) and Art. 78 AIA. If the notified body can confirm the conformity of the high-risk system, a certificate - similar to the audit statement under the DSA - is issued. [^62] Art. 44(1) and (2) AIA. This certificate is valid for up to five years or is revoked or restricted if the AI system subsequently fails to comply with the requirements of the AIA. [^63] Art. 44(3) AIA In order to adapt the legal framework, the Commission is empowered to adopt instruments amending the provisions on conformity assessment procedures. [^64] Art. 43(6) and Art. 97, Rec. (173) AIA.